# Nicolás Honorato Droguett 11    Kazuhiro Kurita 11 $0$-0002-7638-3322    Tesshu Hanaka 22 $0$-0001-6943-856X    Hirotaka Ono 11 $0$-0003-0845-3947

### 28 Dec 2023 v1

@@ Abstract
We propose a new model for graph editing problems on intersection graphs. In well-studied graph editing problems, adding and deleting vertices and edges are used as graph editing operations. As a graph editing operation on intersection graphs, we propose moving objects corresponding to vertices. In this paper, we focus on interval graphs as an intersection graph. We give a linear-time algorithm to find the total moving distance for transforming an interval graph into a complete graph. The concept of this algorithm can be applied for (i) transforming a unit square graph into a complete graph over $L_{\infty}$ distance and (ii) attaining the existence of a $k$-clique on unit interval graphs. In addition, we provide LP-formulations to achieve several properties in the associated graph of unit intervals.
@@ 1 Introduction
The graph editing problem is a problem that consists of editing a graph $G$ so that $G$ satisfies a graph property $\Pi$. Common graph editing operations in well-studied graph editing problems are vertex and edge deletion/addition. The minimum number of such operations required for $G$ to satisfy $\Pi$ is called graph edit distance (GED) [Riesen2015]. When $\Pi$ contains only the target graph, our problem is equivalent to the problem addressed in this paper [Riesen2015]. GED is a metric used to measure the dissimilarity between two given graphs, defined by Sanfeliu and Fu [Sanfeliu1983]. The general problem of determining the GED is known to be NP-hard [Zeng2009]. However, solutions to this problem continue to be designed due to the growing interest in GED as a pairwise graph similarity measure in a wide range of fields such as pattern recognition, image analysis, and molecular analysis [Gao2009, Xiao2013, GarciaHernandez2019]. For instance, in a face recognition scenario, if a facial pattern can be modelled as a graph, and an image is also represented as a graph (called the data graph), then GED can be employed on these graphs to quantify the similarity of the image and the facial model.
In this paper, we propose a new graph edit operation for the graph edit problem on intersection graphs. Given a collection of geometric objects, an intersection graph $G=(V,E)$ is a graph such that $V$ is a collection of geometric objects and $E$ has an edge if two geometric objects intersect. In this graph model, one natural graph edit operation is moving a geometric object. We consider this moving operation as a graph editing operation and address the problem of measuring the similarity between a given graph and a specific graph property. We formulate the problem as follows.
@@
Table 1: Summary of results. In this table, IG, UIG, and USG ($L_{\infty}$) are abbreviations of interval graphs, unit interval graphs, and unit square graphs for $L_{\infty}$ distance, respectively.
@@
As a type of problem that minimises the maximum moving distance, Fomin et al. [fomin2023] addressed the problem of disk dispersal, where a family $\mathcal{S}$ of $n$ disks, an integer $k\geq 0$ and a real $d\geq 0$ are given, and the goal is to decide whether it is possible to obtain a disjoint arrangement by moving at most $k$ disks by at most $d$ distance each. According to the results of Fiala et al. [Fiala2005], this problem is NP-hard when $d=2$ and $k=n$. On the other hand, it was shown by Fomin et al.[fomin2023] that it is FPT when parameterised by $k+d$ and W[1]-hard when parameterised by $k$ if the movement of disks is restricted to rectilinear directions. While some of the cases presented in these works remain open, the system of distant representatives is continuously addressed in the literature [Dumitrescu2015, biedl2021].
This paper is organised as follows. In Section 2, we formally describe the definitions needed to address the above ideas. In Section 3, over Euclidean distance, we show that Problem 1 is solvable in (i) $O(n)$ time for satisfying the property $\Pi_{\texttt{comp}}$ given an interval graph and (ii) $O(n)$ time for unit square graphs for satisfying $\Pi_{\texttt{comp}}$ under standard word-RAM model. In Section 5, we give LP-formulations for satisfying several fundamental graph properties. Lastly, we present our concluding remarks in Section 6.
@@ 2 Preliminaries
In this section, we give general definitions that are used throughout the paper. We refer to the basic terminology of graphs, linear programming, and convex functions described in the textbooks [Diestel:BOOK:2010, cormen2009, boyd2004].
@@
A unit square is a regular quadrilateral where each side has length $1$, and its centre is positioned at $p_{i}=(x_{i},y_{i})$.
@@ 3 Linear-time Algorithms for Satisfying $\Pi_{\texttt{comp}}$
This section gives two liner-time algorithms for satisfying $\Pi_{\texttt{comp}}$ on interval graphs and unit square graphs over Euclidean and $L_{\infty}$ distances, respectively. As a remarkable fact, our algorithm runs even if the given collections of intervals and unit squares are not sorted. We first give a linear-time algorithm for interval graphs.
@@ 3.1 Interval Graphs
First, observe that if an interval graph is complete, then all the intervals share at least one point. We call such a point as a gathering point. Furthermore, if the value of a gathering point $x$ is known and the intervals are scattered over the line, then intervals in $L(\mathcal{I},x)$ have to be moved to the right, and intervals $R(\mathcal{I},x)$ have to be moved to the left. We try to find an optimal $x$, which is called an optimal gathering point. We show that $E(\mathcal{I})$ contains at least one optimal gathering point. We define a function to calculate the moving distance for an arbitrary interval. In particular, the moving distance function is defined as:
@@ 3.2 Unit Square Graphs
We show that Problem 1 applied to unit square graphs is linear-time solvable for $\Pi_{\texttt{comp}}$. Recall that over $L_{\infty}$ distance, the distance between two points $p_{1}=(x_{1},y_{1})$, $p_{2}=(x_{2},y_{2})$ in the plane is: $D(p_{1},p_{2})=\max(|x_{2}-x_{1}|,|y_{2}-y_{1}|)$.
@@
Given a collection of unit squares $\mathcal{S}$, an optimal gathering point can be found in $O(n)$ time for satisfying $\Pi_{\texttt{comp}}$ over $L_{\infty}$ distance.
@@
Observe that the distance calculation depends on the difference between the $x$- and $y$-axis values individually. Moreover, when employing this specific distance metric, two squares intersect if the distance between their centres is less than 1. Hence it is necessary to ensure that the maximum distance between the $x$- and $y$-axis values remains within 1.
To find an optimal gathering point, we utilise Algorithm 1 and apply it separately to each axis. If $p_{x}$ and $p_{y}$ are the optimal gathering points given by the algorithm applied to $x$- and $y$-axis, respectively, then point $(p_{x},p_{y})$ is an optimal gathering point for this case. This can be done in $O(n)$ time. $\square$
@@ 4 Satisfying $\Pi_{k\texttt{-clique}}$ on Unit Interval Graphs in $O(n\log n)$ time
The previous lemma indicates that examining sequences of $k$ consecutive intervals of $\mathcal{I}$ is sufficient to obtain a $k$-clique using the minimum moving distance. Furthermore, using Theorem 3.2, we know that the $k$- and $k+1$-th endpoints are optimal gathering points. Thus we only need to check these endpoints on each sequence to obtain an optimal solution.
The outline of our algorithm is as follows. Let $x_{1}$ and $x_{2}$ be the optimal gathering points of $I(1,k)$ and $I(2,k)$. From Theorem 3.2, $x_{1}$ and $x_{2}$ are medians of $I(1,k)$ and $I(2,k)$. First, we find $x_{1}$ and calculate $D(I(1,k),x_{1})$. Since the difference between $I(1,k)$ and $I(2,k)$ is two, we can find $x_{2}$ and calculate $D(I(2,k),x_{2})$ efficiently. For this purpose, we give formulas for updating the total moving distance.
@@ 1 Procedure $\mathcal{I}$,$k$)
Lemmas 8 and 7 give how to update the total moving distance. To update the total moving distance efficiently based on Lemmas 8, 7 and 5, we need to calculate point $x_{j}^{\mathcal{I}(i,k)}$ for any $j$ and the number of left/right intervals of a given point $x$. It can be done using a binary search tree [cormen2009] and we can show the following theorem.
@@ 5 LP-formulations for Other Fundamental Graph Properties on Unit Interval Graphs
This section deal with the remaining properties, $\Pi_{\texttt{edgeless}}$, $\Pi_{\texttt{acyc}}$, $\overline{\Pi}_{k\texttt{-clique}}$, and $\Pi_{k\texttt{-conn}}$ on unit interval graphs. We show that the problems for satisfying these properties can be formulated using linear programming. It implies that these problems can be solved using polynomial-time algorithms for linear programming.
@@ 5.1 The case $\Pi=\Pi_{\texttt{edgeless}}$
Let $I$ and $J$ be two intervals in $\mathcal{I}$. Without loss of generality, assume that $c(I)\leq c(J)$. Suppose that $c(I)+d_{i}\geq c(J)+d_{j}$. Since $c(I)\leq c(J)$, $d_{i}\geq d_{j}$. In such case, by replacing $d_{i}$ and $d_{j}$ with $c(J)-c(I)+d_{j}$ and $c(I)-c(J)+d_{i}$, we obtain $c(I)+(c(J)-c(I)+d_{j})=c(J)+d_{j}$ and $c(J)+(c(I)-c(J)+d_{i})=c(I)+d_{i}$, respectively. Hence the collection of intervals is the same.
@@
Thus $\mathcal{I}$ has no pair of intervals that intersect when the following inequality holds: $c(I_{i+1})-c(I_{i})>1\text{ for }\leq i\leq n-1$. In other words, the distance between $c(I_{i+1})$ and $c(I_{i})$ must be greater than $1$. Observe that the calculation of the distance does not require applying the absolute function, as $c(I_{i+1})\geq c(I_{i})$. It follows that the following linear program with $O(n)$ constraints can be defined:
@@ 5.2 The case $\Pi=\Pi_{\texttt{acyc}}$
We assume =$\mathcal{I}=\{I_{1},\ldots,I_{n}\}$ is sorted by its centre. Moreover, by Lemma 9, we assume that the order of $\mathcal{I}$ does not change. Since an interval graph is chordal [Jungnickel2013, Thm 9.2.8], we can take any 3-tuple from $\mathcal{I}$ of the form $(I_{i-1},I_{i},I_{i+1})$. This tuple does not make a cycle if at least one pair of intervals is separated by a distance greater than 1. Thus the following condition assures an acyclic unit interval graph: $c(I_{i+2})-c(I_{i})>1$ for $21\leq i\leq n-$ 2. The absolute value can be removed. The following linear program with $O(n)$ constraints is defined:
@@ 5.4 The case $\Pi=\Pi_{k\texttt{-conn}}$
$\begin{array}[]{llcl@{}l}\text{minimise}&\displaystyle\sum\limits_{i=1}^{n}|x_{i}|&\\ \text{subject to}&(c_{i+k+1}+x_{i+k+1})-(c_{i}+x_{i})&\leq&1,&\\ \end{array}$
@@ 6 Concluding Remarks
First, we showed a linear-time algorithm to achieve graph completeness on interval graphs, over Euclidean distance. We demonstrated that graph completeness can be achieved in linear time on unit square graphs over $L_{\infty}$ distance by applying this algorithm. We also depicted the contrast for cases where the moving distance is unique, showing that the optimal gathering points are always the $n$-th and $n+1$-th endpoints in those cases. Using this property, we presented a $O(n\log n)$ running-time algorithm for achieving a $k$-clique on unit interval graphs over Euclidean distance. In addition, we deepen into the property of unit interval graphs used to describe $O(n)$-size linear programs to achieve distinct graph properties on unit interval graphs.

### 09 Jan 2024 v2

@@ Abstract
We propose a new model for graph editing problems on intersection graphs. In well-studied graph editing problems, adding and deleting vertices and edges are used as graph editing operations. As a graph editing operation on intersection graphs, we propose moving objects corresponding to vertices. In this paper, we focus on interval graphs as an intersection graph. We give a linear-time algorithm to find the total moving distance for transforming an interval graph into a complete graph. The concept of this algorithm can be applied for (i) transforming a unit square graph into a complete graph over $L_{1}$ distance and (ii) attaining the existence of a $k$-clique on unit interval graphs. In addition, we provide LP-formulations to achieve several properties in the associated graph of unit intervals.
@@ 1 Introduction
The graph editing problem is a problem of editing a graph $G$ so that $G$ satisfies a graph property $\Pi$. Common graph editing operations in well-studied graph editing problems are vertex and edge deletion/addition. The minimum number of such operations required for $G$ to satisfy $\Pi$ is called graph edit distance (GED) [Riesen2015]. When $\Pi$ contains only the target graph, our problem is equivalent to the problem addressed in [Riesen2015]. GED is a metric used to measure the dissimilarity between two given graphs, defined by Sanfeliu and Fu [Sanfeliu1983]. The general problem of determining the GED is known to be NP-hard [Zeng2009]. However, solutions to this problem continue to be designed due to the growing interest in GED as a pairwise graph similarity measure in a wide range of fields such as pattern recognition, image analysis, and molecular analysis [Gao2009, GarciaHernandez2019, Xiao2013]. For instance, in a face recognition scenario, if a facial pattern can be modelled as a graph, and an image is also represented as a graph, then GED can be employed on these graphs to quantify the similarity of the image and the facial model.
In this paper, we propose a new graph edit operation for the graph edit problem on intersection graphs. Given a collection of geometric objects, an intersection graph $G=(V,E)$ is a graph such that $V$ is a collection of geometric objects represented by vertices and $E$ has an edge if two geometric objects intersect. In this graph model, one natural graph edit operation is moving a geometric object. We consider this moving operation as a graph editing operation and address the problem of measuring the similarity between a given graph and a specific graph property. We formulate the problem as follows.
@@
Table 1: Summary of our results. In this table, IG, UIG, and USG ($L_{1}$) are abbreviations of interval graphs, unit interval graphs, and unit square graphs for $L_{1}$ distance, respectively.
@@
As another type of problem that minimises the maximum moving distance, Fomin et al. [fomin2023] addressed the problem of disk dispersal, where a family $\mathcal{S}$ of $n$ disks, an integer $k\geq 0$ and a real $d\geq 0$ are given, and the goal is to decide whether it is possible to obtain a disjoint arrangement by moving at most $k$ disks by at most $d$ distance each. According to the results of Fiala et al. [Fiala2005], this problem is NP-hard when $d=2$ and $k=n$. On the other hand, it was shown by Fomin et al.[fomin2023] that it is FPT when parameterised by $k+d$ and W[1]-hard when parameterised by $k$ when the movement of disks is restricted to rectilinear directions. While some of the cases presented in these works remain open, the system of distant representatives is continuously addressed in the literature [biedl2021, Dumitrescu2015].
This paper is organised as follows. In Section 2, we formally describe the definitions needed to address the above ideas. In Section 3, we show that Problem 1 is solvable in (i) $O(n)$ time for satisfying the property $\Pi_{\texttt{comp}}$ given an interval graph over Euclidean distance and (ii) $O(n)$ time for unit square graphs for satisfying $\Pi_{\texttt{comp}}$ over $L_{1}$ distance. We use the standard word-RAM model. In Section 4, we present an $O(n\log n)$-time algorithm for satisfying $\Pi_{k\texttt{-clique}}$ given a unit interval graph over Euclidean distance. In Section 5, we give LP-formulations for satisfying several fundamental graph properties. Lastly, we present our concluding remarks in Section 6.
@@ 2 Preliminaries
In this section, we give general definitions that are used throughout the paper. We refer to the basic terminology of graphs, linear programming, and convex functions described in the textbooks [boyd2004, cormen2009, Diestel:BOOK:2010].
@@
A unit square $S$ is a regular quadrilateral where each side has length $1$, and its centre is positioned at $(s_{x},s_{y})\in\mathbb{R}^{2}$.
@@ 3 Linear-time Algorithms for Satisfying $\Pi_{\texttt{comp}}$
This section presents two liner-time algorithms for satisfying $\Pi_{\texttt{comp}}$ on interval graphs and unit square graphs over Euclidean and $L_{1}$ distances, respectively. We remark that our algorithm runs even if the given collections of intervals and unit squares are not sorted. We first give a linear-time algorithm for interval graphs.
@@ 3.1 Interval Graphs
Observe that if an interval graph is complete, then all the intervals share at least one point. We call such a point as a gathering point. Furthermore, if the value of a gathering point $x$ is known and the intervals are scattered over the line, then intervals in $L(\mathcal{I},x)$ have to be moved to the right, and intervals $R(\mathcal{I},x)$ have to be moved to the left. We try to find an optimal $x$, which is called an optimal gathering point. We show that $E(\mathcal{I})$ contains at least one optimal gathering point. We define a function to calculate the moving distance for an arbitrary interval. In particular, the moving distance function is defined as:
@@ 3.2 Unit Square Graphs
We show that Problem 1 applied to unit square graphs is linear-time solvable for $\Pi_{\texttt{comp}}$. Recall that over $L_{1}$ distance, the distance between two points $p_{1}=(x_{1},y_{1})$, $p_{2}=(x_{2},y_{2})$ in the plane is $|x_{2}-x_{1}|+|y_{2}-y_{1}|$. For each unit square $S$, we define the moving distance function $d_{S}$ as follows:
Definition 2.
The moving distance of an unit square $S\in\mathcal{S}$ with centre $(s_{x},s_{y})$ to a point $p=(p_{x},p_{y})$ is a function $d_{S}:\mathbb{R}^{2}\rightarrow\mathbb{R}$ defined as:
$d_{S}(p)=m_{S}(\max(|s_{x}-p_{x}|-1/2,0)+\max(|s_{y}-p_{y}|-1/2,0))$
where $m_{S}\in\mathbb{R}^{+}$ is an arbitrary constant describing the slope. The total moving distance of a collection of squares is then defined as $D(\mathcal{S},p)=\sum_{S\in\mathcal{S}}d_{S}(p)$.
For a given collection of unit squares $\mathcal{S}$, we give a linear-time algorithm that finds a point $p$ that minimises the total moving distance function. Similarly as of intervals, the point $p$ is an optimal gathering point if $p$ minimises the total moving distance function.
@@
Given a collection of unit squares $\mathcal{S}$, an optimal gathering point can be found in $O(n)$ time for satisfying $\Pi_{\texttt{comp}}$ over $L_{1}$ distance.
@@
Let $\mathcal{S}=\{S_{1},\ldots,S_{n}\}$ be a collection of unit squares. We denote the centre of $S_{i}$ as $(s^{i}_{x},s^{i}_{y})$. For $\mathcal{S}$, we consider the collection of unit intervals $\mathcal{I}_{x}=\{I_{1},\ldots,I_{n}\}$ such that the centre of each $I_{i}$ is $s^{i}_{x}$. Similarly, we define the collection of unit intervals $\mathcal{I}_{y}$. Let $p_{x}$ and $p_{y}$ be two optimal gathering points of $\mathcal{I}_{x}$ and $\mathcal{I}_{y}$, respectively. We show that the point $p=(p_{x},p_{y})$ minimises the sum of moving distances for $\mathcal{S}$.
Suppose that $p$ is not an optimal gathering point of $\mathcal{S}$. Let $\mathrm{opt}=(\mathrm{opt}_{x},\mathrm{opt}_{y})$ be an optimal gathering point of $\mathcal{S}$. Since $D(\mathcal{S},p)>D(\mathcal{S},\mathrm{opt})$, $\sum_{I\in\mathcal{I}_{x}}d_{I}(p_{x})>\sum_{I_{i}\in\mathcal{I}_{x}}d_{I}(\mathrm{opt}_{x})$ or $\sum_{I\in\mathcal{I}_{y}}d_{I}(p_{y})>\sum_{I_{i}\in\mathcal{I}_{y}}d_{I}(\mathrm{opt}_{y})$ hold. However, from the definition of $p_{x}$ and $p_{y}$, it contradicts the optimality of $p_{x}$ and $p_{y}$. $\square$
@@ 4 Satisfying $\Pi_{k\texttt{-clique}}$ on Unit Interval Graphs in $O(n\log n)$ time
The previous lemma indicates that examining sequences of $k$ consecutive intervals of $\mathcal{I}$ is sufficient to obtain a $k$-clique using the minimum moving distance. Furthermore, we deduce from Theorem 3.2 that the $k$- and $k+1$-th endpoints are optimal gathering points. Thus we only need to check these endpoints on each sequence to obtain an optimal solution.
The outline of our algorithm is as follows. Let $x_{1}$ and $x_{2}$ be the optimal gathering points of $\mathcal{I}(1,k)$ and $\mathcal{I}(2,k)$. From Theorem 3.2, $x_{1}$ and $x_{2}$ are medians of $\mathcal{I}(1,k)$ and $\mathcal{I}(2,k)$. First, we find $x_{1}$ and calculate $D(\mathcal{I}(1,k),x_{1})$. Since the difference between $\mathcal{I}(1,k)$ and $\mathcal{I}(2,k)$ is two intervals, we can find $x_{2}$ and calculate $D(\mathcal{I}(2,k),x_{2})$ efficiently. For this purpose, we give formulas for updating the total moving distance.
@@ 1 Procedure $\mathcal{I}$,$k$)
Lemmas 8 and 7 provide a method to update the total moving distance. To update the total moving distance efficiently based on Lemmas 8, 7 and 5, we need to calculate point $x_{j}^{\mathcal{I}(i,k)}$ for any $j$ and the number of left/right intervals of a given point $x$. It can be done using a binary search tree [cormen2009] and we can show the following theorem.
@@ 5 LP-formulations for Other Fundamental Graph Properties on Unit Interval Graphs
This section addresses the remaining properties, $\Pi_{\texttt{edgeless}}$, $\Pi_{\texttt{acyc}}$, $\overline{\Pi}_{k\texttt{-clique}}$, and $\Pi_{k\texttt{-conn}}$ on unit interval graphs. We show that the problems for satisfying these properties can be formulated using linear programming. It implies that these problems can be solved by using polynomial-time algorithms for linear programming.
@@ 5.1 The case $\Pi=\Pi_{\texttt{edgeless}}$
Let $I$ and $J$ be two intervals in $\mathcal{I}$. Without loss of generality, assume that $c(I)\leq c(J)$. Suppose that $c(I)+d_{i}\geq c(J)+d_{j}$. Since $c(I)\leq c(J)$, $d_{i}\geq d_{j}$. In this case, by replacing $d_{i}$ and $d_{j}$ with $c(J)-c(I)+d_{j}$ and $c(I)-c(J)+d_{i}$, we obtain $c(I)+(c(J)-c(I)+d_{j})=c(J)+d_{j}$ and $c(J)+(c(I)-c(J)+d_{i})=c(I)+d_{i}$, respectively. Hence the collection of intervals is the same.
@@
Lemma 9 implies that $\mathcal{I}$ has no pair of intervals that intersect when the following inequality holds: $(c(I_{i+1})+d_{i+1})-(c(I_{i})+d_{i})>1\text{ for }\leq i\leq n-1$. It follows that the following linear program with $O(n)$ constraints can be defined:
@@ 5.2 The case $\Pi=\Pi_{\texttt{acyc}}$
We assume =$\mathcal{I}=\{I_{1},\ldots,I_{n}\}$ is sorted by its centre. Moreover, from Lemma 9 we deduce that the order of $\mathcal{I}$ does not change. Since an interval graph is chordal [Jungnickel2013, Thm 9.2.8], we can take any 3-tuple from $\mathcal{I}$ of the form $(I_{i-1},I_{i},I_{i+1})$. This tuple does not make a cycle if at least one pair of intervals does not intersect. Thus the condition $c(I_{i+2})-c(I_{i})>1$ for $21\leq i\leq n-$ 2, assures an acyclic unit interval graph. The following linear program with $O(n)$ constraints is defined:
@@ 5.4 The case $\Pi=\Pi_{k\texttt{-conn}}$
$\begin{array}[]{llcl@{}l}\text{minimise}&\displaystyle\sum\limits_{i=1}^{n}|x_{i}|&\\ \text{subject to}&(c_{i+k+1}+x_{i+k+1})-(c_{i}+x_{i})&\leq&1,&\end{array}$
@@ 6 Concluding Remarks
First, we showed a linear-time algorithm to achieve graph completeness on interval graphs using the minimum moving distance, over Euclidean distance. By applying this algorithm, we demonstrated that graph completeness can be achieved in linear time on unit square graphs over $L_{1}$ distance. We also depicted the contrast for cases where the moving distance is unique, showing that in those cases the optimal gathering points are always the $n$-th and $n+1$-th endpoints. Using this property, we presented a $O(n\log n)$-time algorithm for achieving a $k$-clique on unit interval graphs using the minimum moving distance, over Euclidean distance. In addition, we deepen into the properties of unit interval graphs and describe $O(n)$-size linear programs to achieve distinct graph properties on unit interval graphs.